OpenStudy (anonymous):

A painting is purchased for $450. If the value of the painting doubles every 5 years, then its value is given by the function V(t) = 450 • 2t/5, where t is the number of years since it was purchased and V(t) is its value (in dollars) at that time. What is the value of the painting ten years after its purchase? $1,000 $1,400 $1,800 $2,000 can someone do this problem for me would be a big help.

OpenStudy (anonymous):

replace \(t\) by 10 and see what you get

OpenStudy (anonymous):

2250

OpenStudy (anonymous):

or just think " ten years means it doubled twice" so if it starts at $450 in five years it is worth $900 and in ten years it worth $1800

OpenStudy (anonymous):

it is c then?

OpenStudy (anonymous):

\(450\times 2^{\frac{t}{5}}\) is your formula so replace t by 10 get \[450\times 2^{\frac{10}{5}}=450\times 2^2=450\times 4=1800\]

OpenStudy (anonymous):

yes C it is always C

OpenStudy (anonymous):

dang man your hella fast

OpenStudy (anonymous):

always c ftw